OpenStudy (anonymous):

Mr. andi has 1900 ducks and chickens in the farm altogether. 2/5 of the number of chickens is equal to 1/7 of the number of ducks. the difference between the number o ducks and chickens in the farm is.....

OpenStudy (stamp):

1900 = d + c 2/5 c = 1/7 d c = 5/2 (1/7 d) c = 5/14 d Substitute. 1900 = d + ( 5/14 d ) Simplify. 1900 = 19/14 d Solve. 14/19 ( 1900 ) = d 1400 = d Find c. 1900 = c + d 1900 = c + ( 1400 ) c = 500 Check. 2/5 c = 1/7 d 2/5 ( 500 ) = 1/7 ( 1400 ) 200 = 200